Tossing 4 coins simultaneously finds the probability of getting 2 heads and 2 tails. 


1
Expert's answer
2021-05-07T08:53:54-0400

Assuming a fair coin, independent tosses and 0 chance of landing on the edge.

There are 16 possible results. "\binom{4}{2}=4\times\frac{3}{2} = 6" of them have 2 heads. Since all 16 are equally likely, the chance is "\frac{6}{16} = \frac{3}{8} = 0.375"